OpenStudy (anonymous):

Show that \[\sum_{n+1}^{2n}(6r+4)=n(9n+7)\]

OpenStudy (anonymous):

hint\[\sum_{r=n+1}^{2n} (6r+4)=\sum_{r=1}^{2n} (6r+4)-\sum_{r=1}^{n} (6r+4)\] and use this\[\sum_{k=1}^{n} k=\frac{n(n+1)}{2} \]

OpenStudy (anonymous):

Arithmetic Sequence with r=6 The first term is 4 + 6 (1 + n) the last term is 4 + 6 (1 + n) + 6 (n - 1) \[ S =\frac n 2 ( a_1 + a_n)= \\ S =\frac n 2 (4 + 6 (1 + n) +4 + 6 (1 + n) + 6 (n - 1))=n (7 + 9 n) \]

OpenStudy (anonymous):

lol...much better and simpler...thank u professor

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@mukushla you do you subtract??